OpenStudy (anonymous):

How do you solve 2[4(9-7)+1] please explain the steps?

OpenStudy (anonymous):

Ok to solve this

OpenStudy (anonymous):

Do you remember PEMDAS?

OpenStudy (anonymous):

Yes I do but, it is confusing.

OpenStudy (anonymous):

Don't worry I'll help

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Ok the answer is 18, but I will explain first

OpenStudy (anonymous):

2[4(9-7)+1]

OpenStudy (anonymous):

9-7=2

OpenStudy (anonymous):

2[4(2)+1]

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

4*2= 8

OpenStudy (anonymous):

+1

OpenStudy (anonymous):

you get 2[9]

OpenStudy (anonymous):

Which gives you 18

OpenStudy (anonymous):

If you have another question let me know

OpenStudy (anonymous):

Thank you so much for your help. I understand it now.

OpenStudy (anonymous):

Ok do you have any other questions?

OpenStudy (anonymous):

Just follow the rule PEMDAS

OpenStudy (anonymous):

No I don't have any more questions and ok I will remember to just follow PEMDAS.

OpenStudy (anonymous):

Ok, dont forget :)

OpenStudy (anonymous):

I won't :)

OpenStudy (texaschic101):

These problems can sometimes be confusing, but just take it one step at a time and just like Microbot said, follow pemdas

OpenStudy (anonymous):

if you remember BODMAS rule B-bracket O-order, square roots and powers D-division M-multiplication A-addition S-subtraction 2[4(9-7)+1] first we will solve brackets according to rule =2[4(2)+1] =2[8+1] =16+2 now we use addition 16+2=18

OpenStudy (anonymous):

Ok thank you for all your help.